Mable noticed that 60% of the books in the library were fiction books. She concluded that the ration of fiction to non fiction books was 3:5. Correct or incorrect?
Found 2 solutions by addingup, josgarithmetic:
Answer by addingup(3677)

We divide the % by 100: 60/100
Reduce the fraction to its lowest term: 60/100 = 6/10 = 3/5 which we can also write as 3:5

Answer by josgarithmetic(39617)

If 60 books fiction, and if the rest nonfiction, then ratio of fiction to nonfiction would be 6:4 or 3:2.